# Anchored Supervised Fine-Tuning

### v2

@@ Abstract
Post-training of large language models involves a fundamental trade-off between supervised fine-tuning (SFT), which efficiently mimics demonstrations but tends to memorize, and reinforcement learning (RL), which achieves better generalization at higher computational cost. Dynamic Fine-Tuning (DFT) recently emerged as a promising middle ground, reweighting SFT objectives with token probabilities and achieving improvements in certain reasoning domains, though it exhibits instability in other tasks. We provide a analysis of DFT through the reward-weighted regression (RWR) framework, revealing that it corresponds to a specific auxiliary distribution choice that yields provably tighter RL bounds than standard SFT. However, our analysis also uncovers a critical limitation: this construction lacks distributional anchoring, leading to progressive drift that undermines training stability. To address this, we propose Anchored Supervised Fine-Tuning (ASFT), which augments DFT’s reweighting with lightweight KL regularization to preserve tightness while ensuring stability. Empirically, ASFT consistently outperforms both SFT and DFT across mathematical reasoning, medical knowledge grounding, and code generation, achieving substantial improvements with minimal computational overhead. Our RWR framework provides a systematic lens for understanding post-training methods and demonstrates that principled theoretical analysis leads to both stronger guarantees and practical gains.
11footnotetext: The code will be available at https://github.com/zhuchichi56/ASFT
@@ 1 Introduction
To address these problems, we provide a principled theoretical analysis of DFT within a reward-weighted regression framework inspired by reward-weighted regression and importance sampling theory (Rubinstein & Kroese, 2016). From the theoretical perspective, we reveal that DFT corresponds to a specific auxiliary distribution construction that yields a provably tighter lower bound on the RL objective compared to standard SFT. This causes a critical limitation: the absence of distributional anchoring mechanisms leads to progressive drift away from the reference distribution. The distributional shift undermines the validity of the reward-weighted framework and explains the training instabilities of DFT.
@@
Our contributions are threefold: (1) We provide a rigorous theoretical explanation for DFT’s domain-specific effectiveness and inherent instabilities, grounding its heuristic design within the formal reward-weighted regression framework and proving that it achieves a strictly tighter bound than SFT while suffering from uncontrolled variance growth. (2) We propose ASFT, a simple yet principled method that resolves DFT’s stability issues through lightweight KL anchoring while maintaining its tightness advantages, requiring minimal computational overhead compared to full RL approaches. (3) We demonstrate that ASFT delivers superior performance across both reasoning-intensive and knowledge-intensive domains, achieving better generalization than SFT, greater stability and broader applicability than DFT, and RL-comparable performance with SFT-level computational efficiency.
@@ 2 Related Work
Supervised fine-tuning (SFT) and reinforcement learning (RL) are the two dominant paradigms for post-training large language models. SFT can be viewed as optimizing a stable but loose lower bound on the RL objective, which ensures efficiency and robustness but often leads to memorization and limited generalization (Wei et al., 2022; Chung et al., 2024; Zhang et al., 2021; Chu et al., 2025). In contrast, RL directly optimizes outcome-based rewards, achieving tighter alignment and stronger generalization but at the cost of instability, high variance, and heavy computation (Ouyang et al., 2022; Schulman et al., 2015; 2017). Recent work seeks to bridge this trade-off by either tightening the SFT bound via reweighting and importance weighting (Wu et al., 2025a; Qin & Springenberg, 2025) or stabilizing RL through trust-region and hybrid methods (Sheng et al., 2025; Zhu et al., 2025).
@@
The connection between supervised learning and reinforcement learning through importance weighting has deep theoretical roots (Kahn & Marshall, 1953; Dayan & Hinton, 1997). In the context of policy optimization, importance sampling enables off-policy learning by reweighting samples from a behavior policy to estimate gradients for a target policy (Metelli et al., 2018; Jiang & Li, 2016), though the resulting weights can suffer from high variance when distributions are misaligned (Andradóttir et al., 1995). To address this, trust-region methods (Schulman et al., 2015) and proximal policy optimization (Schulman et al., 2017) constrain policy updates to remain close to a reference policy. Building on these ideas, recent advances in language model fine-tuning have explored importance-weighted supervised objectives (Qin & Springenberg, 2025), proximal supervised fine-tuning (Zhu et al., 2025), and probability-based reweighting in dynamic fine-tuning (Wu et al., 2025a), as well as unified frameworks that combine SFT and RL principles (Lv et al., 2025; Wu et al., 2025b). Our work follows this line of research by anchoring the auxiliary distribution itself to the base model, thereby extending importance-weighted approaches with a mechanism that stabilizes training while retaining their theoretical tightness.
@@ 3.2 The Reward-Weighted Regression Framework
Building on prior work in reward-weighted regression and importance sampling (Peters & Schaal, 2007; Rubinstein & Kroese, 2016; Qin & Springenberg, 2025), we adopt the reward-weighted regression (RWR) framework for language model fine-tuning. This framework provides a principled connection between SFT and RL objectives by leveraging importance sampling and auxiliary distributions to construct tighter bounds on the RL objective.
@@ 4.1 Theoretical Analysis of DFT within the RWR Framework
Key Finding 2: DFT achieves provably tighter bounds than SFT. We proved that this auxiliary distribution yields a strictly tighter lower bound on the RL objective compared to standard SFT whenever the policy assigns non-uniform probabilities to demonstrations (detailed proof in Appendix E).
@@
Key Finding 3: DFT suffers from distributional drift. However, our analysis also revealed a critical limitation: the policy distribution progressively diverges from the reference distribution during training. As optimization proceeds, $q$ becomes increasingly concentrated on trajectories with high $p_{\theta}(\tau)$, creating a feedback loop where the model focuses on a diminishing subset of training data. This distributional drift undermines the foundational assumptions of the RWR framework and threatens the validity of the lower bound guarantees.
@@ 4.2 Anchored Supervised Fine-Tuning (ASFT)
$\mathcal{L}_{\mathrm{ASFT}}(\theta)=\mathcal{L}_{\mathrm{DFT}}(\theta)+\lambda\mathbb{E}_{s}[D_{\mathrm{KL}}(\pi_{\theta}(\cdot\mid s)\|\pi_{\mathrm{base}}(\cdot\mid s))]$ (6)
@@ 5.2 Main Results
Model MedQA MMLU MedMCQA Avg AIME24 Math500 Minerva Olympiad ACM23 Avg
LLaMA-2-7B/Qwen2.5-7B 29.85 30.52 33.76 31.38 1.65 28.79 9.26 7.69 15.65 12.61
@@
SFT 33.31 (+3.46) 33.52 (+3.00) 33.28 (-0.48) 33.37 (+1.99) 1.24 (-0.41) 41.84 (+13.05) 11.30 (+2.04) 12.26 (+4.57) 17.03 (+1.38) 16.73 (+4.12)
SFT w/ KL 29.22 (-0.63) 30.63 (+0.11) 33.01 (-0.75) 30.95 (-0.43) 0.41 (-1.24) 42.21 (+13.42) 12.05 (+2.79) 12.08 (+4.39) 17.19 (+1.54) 16.79 (+4.18)
DFT 29.69 (-0.16) 26.69 (-3.83) 31.20 (-2.56) 29.19 (-2.19) 4.18 (+2.53) 59.51 (+30.72) 17.10 (+7.84) 24.95 (+17.26) 31.09 (+15.44) 27.77 (+15.16)
[…]
@@
SFT 33.54 (+3.69) 38.48 (+7.96) 36.03 (+2.27) 36.02 (+4.64) 2.71 (+1.06) 44.74 (+15.95) 13.21 (+3.95) 13.44 (+5.75) 21.56 (+5.91) 19.93 (+7.32)
SFT w/ KL 30.56 (+0.71) 29.86 (-0.66) 33.56 (-0.20) 31.33 (-0.05) 2.70 (+1.05) 44.91 (+16.12) 13.03 (+3.77) 13.48 (+5.79) 18.90 (+3.25) 18.60 (+5.99)
DFT 31.26 (+1.41) 33.08 (+2.56) 35.09 (+1.33) 33.14 (+1.76) 3.34 (+1.69) 57.93 (+29.14) 23.28 (+14.02) 25.31 (+17.62) 28.44 (+12.79) 27.66 (+15.05)
[…]
@@
SFT 33.46 (+3.61) 38.01 (+7.49) 35.67 (+1.91) 35.71 (+4.33) 0.83 (-0.82) 47.30 (+18.51) 13.46 (+4.20) 14.16 (+6.47) 20.00 (+4.35) 19.15 (+6.54)
SFT w/ KL 30.09 (+0.24) 31.62 (+1.10) 33.85 (+0.09) 31.85 (+0.47) 1.44 (-0.21) 46.81 (+17.02) 14.13 (+4.87) 13.74 (+6.05) 20.00 (+4.35) 19.22 (+6.61)
DFT 36.61 (+6.76) 41.26 (+10.74) 36.31 (+2.55) 38.06 (+6.68) 6.26 (+4.61) 56.88 (+28.09) 21.18 (+11.92) 22.68 (+14.99) 27.19 (+11.54) 26.04 (+13.43)
[…]
Table 1: Performance comparison of fine-tuning methods on medical benchmarks (left) and math benchmarks (right) under different dataset scales. Bold numbers indicate the best performance in each group, and rows with blue background highlight our ASFT approach. Numbers in (olive) or (red) show improvements or drops relative to the base model.
@@ 5.3 Abaltion Study
Here we follow the RLHF convention: let $Q$ denote the policy model (i.e., the fine-tuned model with parameters $\theta$), and $P$ denote the base model, which is the pretrained model before any fine-tuning. In our approach, we use the reverse KL divergence, $D_{\mathrm{KL}}(Q\,\|\,P)$, to regularize the policy model towards the base model at each step. For comparison, we also experiment with the forward KL divergence, $D_{\mathrm{KL}}(P\,\|\,Q)$. As shown in Figure 2, reverse KL leads to stable convergence and consistent improvements. In contrast, forward KL tends to encourage mode-covering behavior, causing the model to spread probability mass more broadly. We further sweep the regularization coefficient $\lambda$ and find that an optimal range is around $\lambda=0.1$; both excessively small and large values lead to under-anchoring or over-regularization, respectively.
Figure 2: Comparison of reverse KL (RKL) and forward KL divergence regularization effects across different coefficient values ($\lambda$) on MedQA, MMLU, and MedMCQA benchmarks. Performance is measured in accuracy scores, with horizontal dashed lines indicating baseline performance of Base, SFT, and DFT models for reference.
@@
We conduct an ablation study to evaluate the robustness of ASFT with respect to learning rate and batch size, using LLaMA-2-7B and 10k samples from MedCAQA. For learning rate, we examine six values ranging from 5e-6 to 2e-4. ASFT consistently outperforms both SFT and DFT across all configurations, with the best performance observed at intermediate rates (1e-5 and 2e-4). Extreme low (5e-6) or high (2e-4) rates lead to minor drops, highlighting that moderate learning rates are preferable but ASFT remains robust overall. For batch size, we sweep values from 32 to 256. The results show stable performance across the full range, with only small fluctuations. This indicates that batch size is not a sensitive factor for ASFT, and standard settings suffice. Overall, ASFT demonstrates both strong robustness and low sensitivity to key hyperparameters, maintaining a consistent advantage over SFT and DFT in medical QA fine-tuning tasks. The detailed results for learning rate and batch size sweeps are shown in Appendix G.1.
@@ 6.1 Scaling Analysis Across Model Size
We evaluate fine-tuning methods on medical-domain datasets using 10k training samples (Figure see 3, detailed results in Appendix F.1). Our experiments cover both LLaMA-2 (7B and 70B) and Qwen2.5 (7B, 32B, and 72B) models. Across all model sizes, ASFT consistently outperforms Base, SFT, and DFT, and its improvements remain robust as models scale, demonstrating effective and stable adaptation under low-resource settings.
Model MedQA MMLU MedMCQA Average
@@
DFT 24.67 (-5.18) 22.82 (-7.70) 30.43 (-3.33) 25.97 (-5.41)
iw-SFT 28.36 (+1.49) 35.89 (+5.37) 34.88 (+1.12) 33.04 (+1.66)
SFT 33.31 (+3.56) 33.52 (+3.00) 33.28 (-0.48) 33.44 (+2.06)
ASFT 39.28 (+9.43) 46.37 (+15.85) 40.45 (+6.69) 42.03 (+10.65)
GRPO 30.48 (+0.63) 32.46 (+1.94) 34.64 (+0.88) 32.53 (+1.15)
DAPO 39.75 (+9.90) 48.63 (+18.11) 38.37 (+4.61) 42.25 (+10.87)
SFT + DAPO 36.84 (+6.99) 44.76 (+14.24) 39.11 (+5.35) 40.24 (+8.86)
ASFT + DAPO 41.32 (+11.47) 49.54 (+19.02) 41.45 (+7.69) 44.10 (+12.72)
Table 2: Comparison of different post-training strategies on medical benchmarks. Numbers in parentheses indicate differences relative to LLaMA-2-7B.

### v3

@@ Abstract
Post-training of large language models involves a fundamental trade-off between supervised fine-tuning (SFT), which efficiently mimics demonstrations but tends to memorize, and reinforcement learning (RL), which achieves better generalization at higher computational cost. Dynamic Fine-Tuning (DFT) recently emerged as a promising middle ground, reweighting SFT objectives with token probabilities and achieving improvements in certain reasoning domains, though it exhibits instability in other tasks. We provide an analysis of DFT through the reward-weighted regression (RWR) framework, revealing that it corresponds to a specific auxiliary distribution choice that yields provably tighter RL bounds than standard SFT. However, our analysis also uncovers a critical limitation: this construction lacks distributional anchoring, leading to progressive drift that undermines training stability. To address this, we propose Anchored Supervised Fine-Tuning (ASFT), which augments DFT’s reweighting with lightweight KL regularization to preserve tightness while ensuring stability. Empirically, ASFT consistently outperforms both SFT and DFT across mathematical reasoning, medical knowledge grounding, and code generation, achieving substantial improvements with minimal computational overhead. Our RWR framework provides a systematic lens for understanding post-training methods and demonstrates that principled theoretical analysis leads to both stronger guarantees and practical gains. The code is available at https://github.com/zhuchichi56/ASFT.
@@ 1 Introduction
To address these problems, we provide a principled theoretical analysis of DFT within a reward-weighted regression framework inspired by reward-weighted regression and importance sampling theory (Rubinstein and Kroese, 2016). From the theoretical perspective, we reveal that DFT corresponds to a specific auxiliary distribution construction that yields a provably tighter lower bound on the RL objective compared to standard SFT. This causes a critical limitation: the absence of distributional anchoring mechanisms leads to progressive drift away from the reference distribution. The distributional shift makes the bound increasingly loose and increases importance weight variance, explaining the training instabilities of DFT.
@@
Our contributions are threefold: (1) We provide a rigorous theoretical explanation for DFT’s domain-specific effectiveness and inherent instabilities, grounding its heuristic design within the formal reward-weighted regression framework and proving that it achieves a strictly tighter bound than SFT while suffering from uncontrolled variance growth. (2) We propose ASFT, a simple yet principled method that resolves DFT’s stability issues through lightweight KL anchoring while maintaining its tightness advantages, requiring minimal computational overhead compared to full RL approaches. (3) We demonstrate that ASFT delivers superior performance across both reasoning-intensive and knowledge-intensive domains, achieving better generalization than SFT, greater stability and broader applicability than DFT, and RL-comparable performance with SFT-level computational efficiency [^1] The code is available at https://github.com/zhuchichi56/ASFT..
@@ 2 Related Work
Supervised fine-tuning (SFT) and reinforcement learning (RL) are the two dominant paradigms for post-training large language models. SFT can be viewed as optimizing a stable but loose lower bound on the RL objective, which ensures efficiency and robustness but often leads to memorization and limited generalization (Wei et al., 2022; Chung et al., 2024; Zhang et al., 2021; Chu et al., 2025). In contrast, RL directly optimizes outcome-based rewards, achieving tighter alignment and stronger generalization but at the cost of instability, high variance, and heavy computation (Ouyang et al., 2022; Schulman et al., 2015; 2017). Recent work seeks to bridge this trade-off by either tightening the SFT bound via reweighting and importance weighting (Wu et al., 2025a; Qin and Springenberg, 2025) or stabilizing RL through trust-region and hybrid methods (Sheng et al., 2025; Zhu et al., 2025).
@@
The connection between supervised learning and reinforcement learning through importance weighting has deep theoretical roots (Kahn and Marshall, 1953; Dayan and Hinton, 1997). In the context of policy optimization, importance sampling enables off-policy learning by reweighting samples from a behavior policy to estimate gradients for a target policy (Metelli et al., 2018; Jiang and Li, 2016), though the resulting weights can suffer from high variance when distributions are misaligned (Andradóttir et al., 1995). To address this, trust-region methods (Schulman et al., 2015) and proximal policy optimization (Schulman et al., 2017) constrain policy updates to remain close to a reference policy. Building on these ideas, recent advances in language model fine-tuning have explored importance-weighted supervised objectives (Qin and Springenberg, 2025), proximal supervised fine-tuning (Zhu et al., 2025), and probability-based reweighting in dynamic fine-tuning (Wu et al., 2025a), as well as unified frameworks that combine SFT and RL principles (Lv et al., 2025; Wu et al., 2025b). Our work follows this line of research by anchoring the auxiliary distribution itself to the base model, thereby extending importance-weighted approaches with a mechanism that stabilizes training while retaining their theoretical tightness.
@@ 3.2 The Reward-Weighted Regression Framework
Building on prior work in reward-weighted regression and importance sampling (Peters and Schaal, 2007; Rubinstein and Kroese, 2016; Qin and Springenberg, 2025), we adopt the reward-weighted regression (RWR) framework for language model fine-tuning. This framework provides a principled connection between SFT and RL objectives by leveraging importance sampling and auxiliary distributions to construct tighter bounds on the RL objective.
@@ 4.1 Theoretical Analysis of DFT within the RWR Framework
Key Finding 2: DFT achieves provably tighter bounds than SFT. We proved that this auxiliary distribution yields a strictly tighter lower bound on the RL objective compared to standard SFT whenever the policy assigns non-uniform probabilities to demonstrations (detailed proof in Appendix D.4).
@@
Key Finding 3: DFT suffers from distributional drift. However, our analysis also revealed a critical limitation: the policy distribution progressively diverges from the reference distribution during training. As optimization proceeds, $q$ becomes increasingly concentrated on trajectories with high $p_{\theta}(\tau)$, creating a feedback loop where the model focuses on a diminishing subset of training data. This distributional drift makes the bound increasingly loose and the importance weights increasingly high-variance, reducing effective sample size and destabilizing optimization.
@@ 4.2 Anchored Supervised Fine-Tuning (ASFT)
$\mathcal{L}_{\mathrm{ASFT}}(\theta)=\mathcal{L}_{\mathrm{DFT}}(\theta)+\lambda\mathbb{E}_{s}[D_{\mathrm{KL}}(\pi_{\mathrm{base}}(\cdot\mid s)\|\pi_{\theta}(\cdot\mid s))]$ (6)
@@ 5.2 Main Results
Methods Medical Benchmarks Math Benchmarks
MedQA MMLU MedMCQA Avg. AIME24 Math500 Minerva Olympiad ACM23 Avg.
Base 29.85 30.52 33.76 31.38 1.65 28.79 9.26 7.69 15.65 12.61
@@
SFT 33.31 $\uparrow$3.46 33.52 $\uparrow$3.00 33.28 $\downarrow$0.48 33.37 $\uparrow$1.99 1.24 $\downarrow$0.41 41.84 $\uparrow$13.05 11.30 $\uparrow$2.04 12.26 $\uparrow$4.57 17.03 $\uparrow$1.38 16.73 $\uparrow$4.12
SFT w/ KL 29.22 $\downarrow$0.63 30.63 $\uparrow$0.11 33.01 $\downarrow$0.75 30.95 $\downarrow$0.43 0.41 $\downarrow$1.24 42.21 $\uparrow$13.42 12.05 $\uparrow$2.79 12.08 $\uparrow$4.39 17.19 $\uparrow$1.54 16.79 $\uparrow$4.18
DFT 29.69 $\downarrow$0.16 26.69 $\downarrow$3.83 31.20 $\downarrow$2.56 29.19 $\downarrow$2.19 4.18 $\uparrow$2.53 59.51 $\uparrow$30.72 17.10 $\uparrow$7.84 24.95 $\uparrow$17.26 31.09 $\uparrow$15.44 27.77 $\uparrow$15.16
ASFT 39.28 $\uparrow$9.43 46.37 $\uparrow$15.85 40.45 $\uparrow$6.69 42.03 $\uparrow$10.65 3.33 $\uparrow$1.68 59.60 $\uparrow$30.81 19.91 $\uparrow$10.65 24.50 $\uparrow$16.81 36.41 $\uparrow$20.76 28.75 $\uparrow$16.14
@@
SFT 33.54 $\uparrow$3.69 38.48 $\uparrow$7.96 36.03 $\uparrow$2.27 36.02 $\uparrow$4.64 2.71 $\uparrow$1.06 44.74 $\uparrow$15.95 13.21 $\uparrow$3.95 13.44 $\uparrow$5.75 21.56 $\uparrow$5.91 19.93 $\uparrow$7.32
SFT w/ KL 30.56 $\uparrow$0.71 29.86 $\downarrow$0.66 33.56 $\downarrow$0.20 31.33 $\downarrow$0.05 2.70 $\uparrow$1.05 44.91 $\uparrow$16.12 13.03 $\uparrow$3.77 13.48 $\uparrow$5.79 18.90 $\uparrow$3.25 18.60 $\uparrow$5.99
DFT 31.26 $\uparrow$1.41 33.08 $\uparrow$2.56 35.09 $\uparrow$1.33 33.14 $\uparrow$1.76 3.34 $\uparrow$1.69 57.93 $\uparrow$29.14 23.28 $\uparrow$14.02 25.31 $\uparrow$17.62 28.44 $\uparrow$12.79 27.66 $\uparrow$15.05
ASFT 42.03 $\uparrow$12.18 44.94 $\uparrow$14.42 39.06 $\uparrow$5.30 42.01 $\uparrow$10.63 5.81 $\uparrow$4.16 57.03 $\uparrow$28.24 20.61 $\uparrow$11.35 24.44 $\uparrow$16.75 30.00 $\uparrow$14.35 27.18 $\uparrow$14.57
@@
SFT 33.46 $\uparrow$3.61 38.01 $\uparrow$7.49 35.67 $\uparrow$1.91 35.71 $\uparrow$4.33 0.83 $\downarrow$0.82 47.30 $\uparrow$18.51 13.46 $\uparrow$4.20 14.16 $\uparrow$6.47 20.00 $\uparrow$4.35 19.15 $\uparrow$6.54
SFT w/ KL 30.09 $\uparrow$0.24 31.62 $\uparrow$1.10 33.85 $\uparrow$0.09 31.85 $\uparrow$0.47 1.44 $\downarrow$0.21 46.81 $\uparrow$17.02 14.13 $\uparrow$4.87 13.74 $\uparrow$6.05 20.00 $\uparrow$4.35 19.22 $\uparrow$6.61
DFT 36.61 $\uparrow$6.76 41.26 $\uparrow$10.74 36.31 $\uparrow$2.55 38.06 $\uparrow$6.68 6.26 $\uparrow$4.61 56.88 $\uparrow$28.09 21.18 $\uparrow$11.92 22.68 $\uparrow$14.99 27.19 $\uparrow$11.54 26.04 $\uparrow$13.43
ASFT 40.61 $\uparrow$10.76 42.02 $\uparrow$11.50 37.32 $\uparrow$3.56 39.98 $\uparrow$8.60 6.66 $\uparrow$5.01 59.99 $\uparrow$31.20 23.55 $\uparrow$14.29 25.57 $\uparrow$17.88 36.72 $\uparrow$21.07 30.50 $\uparrow$17.89
Table 1: Performance comparison of fine-tuning methods on medical benchmarks (left, base: LLaMA-2-7B) and math benchmarks (right, base: Qwen2.5-7B) under different dataset scales. Bold numbers indicate the best performance in each group, and rows with blue background highlight our ASFT approach. Arrows with $\uparrow$ and $\downarrow$ indicate performance improvements and degradations relative to the base model.
@@ 5.3 Ablation Study
Here we follow the standard notation: let $Q$ denote the policy model (i.e., the fine-tuned model with parameters $\theta$), and $P$ denote the reference model (the pretrained base model). In our approach, we use the forward KL divergence, $D_{\mathrm{KL}}(P\,\|\,Q)$, to regularize the policy model. For comparison, we also experiment with the reverse KL divergence, $D_{\mathrm{KL}}(Q\,\|\,P)$. As shown in Figure 2, forward KL leads to consistent improvements. Theoretically, forward KL encourages mode-covering behavior, preventing the model from collapsing and ensuring it maintains the broad distribution of the base model. In contrast, reverse KL is mode-seeking, which can cause the model to focus excessively on a few high-probability sequences. We further sweep the regularization coefficient $\lambda$ and find that the optimal range is around $\lambda=0.1$; both excessively small and large values lead to under-anchoring or over-regularization, respectively.
Figure 2: Comparison of forward KL (FKL) and reverse KL (RKL) regularization effects across different coefficient values ($\lambda$) on MedQA, MMLU, and MedMCQA benchmarks. Performance is measured in accuracy, with horizontal dashed lines indicating baseline performance of Base, SFT, and DFT models.
@@
We conduct an ablation study to evaluate the robustness of ASFT with respect to learning rate and batch size, using LLaMA-2-7B and 10k samples from MedCAQA. For learning rate, we examine six values ranging from 5e-6 to 2e-4. ASFT consistently outperforms both SFT and DFT across all configurations, with the best performance observed at intermediate rates (1e-5 and 2e-4). Extreme low (5e-6) or high (2e-4) rates lead to minor drops, highlighting that moderate learning rates are preferable but ASFT remains robust overall. For batch size, we sweep values from 32 to 256. The results show stable performance across the full range, with only small fluctuations. This indicates that batch size is not a sensitive factor for ASFT, and standard settings suffice. Overall, ASFT demonstrates both strong robustness and low sensitivity to key hyperparameters, maintaining a consistent advantage over SFT and DFT in medical QA fine-tuning tasks. The detailed results for learning rate and batch size sweeps are shown in Appendix E.2.
@@ 6.1 Scaling Analysis Across Model Size
We evaluate fine-tuning methods on medical-domain datasets using 10k training samples (Figure see 3, detailed results in Appendix E.1). Our experiments cover both LLaMA-2 (7B and 70B) and Qwen2.5 (7B, 32B, and 72B) models. Across all model sizes, ASFT consistently outperforms Base, SFT, and DFT, and its improvements remain robust as models scale, demonstrating effective and stable adaptation under low-resource settings.
Methods MedQA MMLU MedMCQA Avg.
@@
Supervised Fine-Tuning
DFT 24.67 $\downarrow$5.18 22.82 $\downarrow$7.70 30.43 $\downarrow$3.33 25.97 $\downarrow$5.41
iw-SFT 28.36 $\uparrow$1.49 35.89 $\uparrow$5.37 34.88 $\uparrow$1.12 33.04 $\uparrow$1.66
SFT 33.31 $\uparrow$3.56 33.52 $\uparrow$3.00 33.28 $\downarrow$0.48 33.44 $\uparrow$2.06
ASFT 39.28 $\uparrow$9.43 46.37 $\uparrow$15.85 40.45 $\uparrow$6.69 42.03 $\uparrow$10.65
Reinforcement Learning
GRPO 30.48 $\uparrow$0.63 32.46 $\uparrow$1.94 34.64 $\uparrow$0.88 32.53 $\uparrow$1.15
DAPO 39.75 $\uparrow$9.90 48.63 $\uparrow$18.11 38.37 $\uparrow$4.61 42.25 $\uparrow$10.87
Continual RL
SFT + DAPO 36.84 $\uparrow$6.99 44.76 $\uparrow$14.24 39.11 $\uparrow$5.35 40.24 $\uparrow$8.86
ASFT + DAPO 41.32 $\uparrow$11.47 49.54 $\uparrow$19.02 41.45 $\uparrow$7.69 44.10 $\uparrow$12.72
Table 2: Comparison of different post-training strategies on medical benchmarks. Bold numbers indicate the best performance in each group, and rows with blue background highlight our ASFT approach.